TANGENTS AND NORMALS 
find the values of a and b if the slope of the tangent to the curve xy + ax + by = 2 at (1,1) is 2.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

xy + ax + by = 2
Slope of tangent:
xy' + y + a + by' = 0
y’ = 2 at (1, 1)
2 + 1 + a + 2b = 0
a + 2b = -3…....(1)
Also line passing through (1, 1)
1 + a + b = 2
a + b = 1….....(2)
Solving (1) and (2), we have
a = 5, b = -4